OpenStudy (anonymous):

Distributive Property 12+5x-2(3x+5)

OpenStudy (lyrae):

\[c(a+b) = ca + cb \] For example \[3(7 + 4x) = 3 \times 7 + 3 \times 4x = 21 + 12x \]